2/3+n+6=3/4 need to solve but dont know what to do with the fractions ugh
Answer by Alan3354(69443)

2/3+n+6=3/4 need to solve but dont know what to do with the fractions ugh
----------
Step 1, multiply thru by 12
8 + 12n + 72 = 9
No fractions